# SAP-CoPE: Social-Aware Planning using Cooperative Pose Estimation with Infrastructure Sensor Nodes

### v1

@@ Abstract
Autonomous driving systems must operate safely in human-populated indoor environments, where challenges such as limited perception and occlusion sensitivity arise when relying solely on onboard sensors. These factors generate difficulties in the accurate recognition of human intentions and the generation of comfortable, socially aware trajectories. To address these issues, we propose SAP-CoPE, a social-aware planning framework that integrates cooperative infrastructure with a novel 3D human pose estimation method and a model predictive control-based controller. This real-time framework formulates an optimization problem that accounts for uncertainty propagation in the camera projection matrix while ensuring human joint coherence. The proposed method is adaptable to single- or multi-camera configurations and can incorporate sparse LiDAR point-cloud data. To enhance safety and comfort in human environments, we integrate a human personal space field based on human pose into a model predictive controller, enabling the system to navigate while avoiding discomfort zones. Extensive evaluations in both simulated and real-world settings demonstrate the effectiveness of our approach in generating socially aware trajectories for autonomous systems.
@@ I Introduction
Intelligent indoor mobility systems are gaining increasing attention in healthcare, logistics, and the automotive industry [1, 2, 3]. These systems reduce manual transportation tasks, minimizing workplace injuries and improving operational efficiency [4]. Achieving these benefits requires the system to have real-time obstacle detection, human trajectory prediction, and coordinated robot motion for safe and smooth navigation.
While research has enhanced robotic perception ability through high-end onboard sensors, these systems face limitations in perception range and field of view (FOV), especially in crowded indoor spaces. Infrastructure Sensor Nodes (ISNs) offer a solution [5]. Ning et al. proposed ceiling-mounted sensor nodes, which expand the FOV, reduce occlusions, and provide a cost-effective alternative to expensive onboard sensors. A centralized network of ISNs supports multiple robots, improving scalability and reducing costs.
Figure 1: Key challenges: (1) How to reliably detect humans in crowded, occluded scenarios? (2) How to control the robot for human-comfortable motion?
Effective human-robot interaction is essential for both system efficiency and safety [6]. Accurate human intention prediction enables proactive navigation and collision avoidance. To achieve this, We propose a 3D pose estimation approach using probabilistic fusion of image and sparse point cloud data, ensuring real-time, accurate human behavior analysis.
Yang et al. [7] introduced a Model Predictive Control (MPC)-based motion planning algorithm that treats all objects as obstacles, prioritizing only on physical safety. However, psychological safety is also equally important when navigating near humans. Research shows that people maintain social zones in shared spaces [8]. To address this, we integrate a Personal Space concept from psychology into motion planning, enabling robots to respect social behavior, maintain comfortable distances, and move naturally in human-populated environments.
@@
    Real-time Cooperative Perception System: We proposed a cooperative perception framework that leveraging sensor node networks to provide comprehensive environmental awareness. This system accounts for both computation and communication latencies, ensuring a timely, synchronized, and overall understanding of dynamic environments.
@@
    3D Human Pose Detection We develop a novel approach for 3D human pose detection that combines geometry-based techniques and sensor fusion between image data and sparse point clouds. This method enhances detection accuracy.
@@
    MPC-based Motion Planning and Control Scheme: We introduced an integrated motion planning and control framework based on MPC. Our approach simultaneously manages both human interactions and obstacle avoidance by incorporating a Personal Space (PS) model for social comfort and Artificial Potential Fields (APF) for dynamic obstacle navigation.
@@
    Validation through Simulation and Real-world Experiments: The proposed system is extensively verified and validated through simulations and real-world experiments. Results demonstrate its effectiveness in handling human-robot interactions, obstacle avoidance, and motion tracking, enabled by accurate perception.
The structure of this article is organized as follows. Section II presents an overview of the proposed methodology. Section III, details the perception algorithm. Section IV formulates the motion planning and tracking controller. Section V demonstrates the simulation and experimental results generated by the proposed system. Finally, Section VI concludes the study and discusses potential future research directions.
@@ II Related Works
In restricted environments, mobile robots primarily focus on avoiding collisions with objects and are programmed to stop when humans are nearby. However, this strategy is less effective in human-populated areas like hospitals and factories, where interactions with people are unavoidable. This section examines the two key components of these algorithms: perception and planning.
@@ II-A Perception
Cooperative Perception System: Cooperative perception enhances environmental awareness by integrating sensory data from multiple sources, including other robots and infrastructure. This collaborative approach effectively addresses challenges such as occlusions and the limited FOV inherent in individual sensors. Recent advancements highlight the significance of Vehicle-to-Everything (V2X) in autonomous driving domain[9, 10, 11, 12]. Huang et al. provide a comprehensive overview of V2X-based cooperative perception, highlighting its evolution and the critical role of reliable data sharing between robots and infrastructure[10]. However, real-time performance and robustness to communication delays remain challenges. To address these limitations, we propose a real-time cooperative perception system utilizing ceiling-mounted sensors called Infrastructure Sensor Nodes (ISNs). This setup expands the FOV, reduces occlusions in crowded spaces, and ensures timely environmental understanding by mitigating computation and communication latencies, enhancing intelligent indoor mobility systems.
3D Human Pose Detection: Accurately estimating human 3D poses is vital for autonomous systems to interpret and predict human behavior, ensuring safe and socially aware interactions. Single camera-based methods have faced challenges such as depth ambiguity and occlusions[13, 14], thus, the estimation accuracy can be enhanced by leveraging multi-modal sensor data. Zanfir et al. introduced HUM3DIL, a semi-supervised multi-modal approach that embed LiDAR points into pixel-aligned multi-modal features and employing a Transformer architecture for prediction[15]. Similarly, Bauer et al. proposed a weakly supervised method that integrates camera and LiDAR data, enabling accurate 3D human pose estimation without extensive 3D annotations[16], however, it needs dense point cloud to generate high-accuracy pseudo 3D keypoint labels via projecting point cloud to image.
Our work introduces a novel probabilistic 3D human pose estimation method that robustly integrates geometry-based techniques with image data and optionally available sparse point cloud measurements. By leveraging probabilistic fusion and joint coherence. Unlike the previous approaches, our approach reliably provides accurate real-time 3D pose predictions, even in the absence of sparse 3D measurements. This flexibility ensures consistent and precise human behavior understanding, particularly in challenging indoor scenarios characterized by occlusions and low visibility.
@@ II-B Planning
Social-aware Planning: Recent human-aware planning algorithms focus primarily on human interactions while often neglecting static and dynamic obstacles[17, 18, 19, 20]. Among these algorithms, the Social Force Model (SFM) is widely used to generate socially acceptable robot motion. Kamezaki et al. [21] introduced a planning algorithm based on the inducible Social Force Model, while Ratsamme et al. [22] developed the Extended SFM (ESFM) by incorporating the human pose, orientation, etc. to improve the performance of conventional SFM. Although these approaches have shown excellent performance in human-aware planning; however, they do not explicitly incorporate normal obstacles into their planning frameworks. To address this gap, we proposed an MPC-based integrated motion planning and tracking algorithm that simultaneously accounts for both human interactions and obstacle avoidance by leveraging the APF and PS model.
Planning based on Information Types: Human-aware planning algorithms can be broadly categorized into two types: individual state-based information and group-based information. Individual state-based methods use detailed information such as position, pose, orientation, and velocity; for example, Stefanini et al. [23] proposed a novel planning algorithm that incorporates the articulated 3D human poses, semantic labels, and trajectory prediction to describe human behavior. In contrast, group-based methods treat multiple humans as a single entity. Truong et al. [24] opposed a proactive social motion model that accounts socio-spatiotemporal characteristics of human groups, allows the mobile robot to plan a safe trajectory in various situations.
@@
In this paper, we adopt the Personal Space (PS) field proposed by Amaoka et al. [25], which defines a non-viable zone that ensures the psychological safety of humans. This approach incorporates an individual’s position and orientation to construct the PS model into the robot’s motion planning algorithm. While our method is based on individual state information, it achieves performance comparable to group-based approaches, providing a balanced solution for socially aware robot navigation.
III Methodology
The proposed SAP-CoPE method combines real-time cooperative perception with socially aware motion planning and control to enhance autonomous mobility in dynamic indoor environments (Fig. 2). It employs a network of multi-modal ISNs equipped with LiDAR, cameras, and edge computing for local object detection and 3D human pose estimation. A synchronized clock system aligns LiDAR scans with camera operations, ensuring temporally consistent data, which is transmitted to a Central Layer (CL) for global fusion, multi-object tracking, and delay compensation.
The MPC-based controller is designed to generate optimal control commands, enabling the robot to follow a collision-free trajectory while adhering to a predefined reference motion. To enhance navigation, the APF model and PS model are incorporated into the MPC’s objective function. APF ensures effective obstacle avoidance, while the PS model maintains human comfort by respecting personal space boundaries.
The optimal control commands are then transmitted to the Application Layer (AL) for execution. The SAP-CoPE framework is adaptable to various robotic platforms, such as the autonomous cart and autonomous medical bed at the Mechatronic Vehicle Systems Lab, University of Waterloo. In this study, only the autonomous medical bed [26] was used for simulations and experimental validation, and referred as term “robot” through out this paper.
@@ IV-A1 Camera-based 2D Detection
Recent advancements in deep learning have significantly improved camera-based 2D object detection. Notably, YOLOv11 [27] demonstrates superior accuracy and inference speed due to its state-of-the-art backbone and neck architectures. However, the pre-trained model does not generalize well to the top-down sensor node view and is unable to detect the autonomous robot used in our application. To address these limitations, we fine-tuned the YOLOv11 model using a dataset combining COCO [28] and the Indoor Cooperative Infrastructure Dataset (ICID) [5].
@@ IV-A2 LiDAR Processing and Class Association
The LiDAR processing consists of region-of-interest filtering and a hierarchical clustering algorithm to accommodate the scanning pattern. The multi-model class association fuses and aligns the data from both camera and LiDAR modalities to enhance perception accuracy. Further details on the implementation can be found in our previous work [5].
@@
Estimating absolute 3D human pose from a single camera is challenging due to the lack of depth information. In our approach, we first obtain the 2D pixel coordinates of each joint and leverage prior height distribution information to estimate the rough 3D distribution of each joint as $P^{i}\sim N(\mu_{P^{i}},\Sigma_{P^{i}})$ in the absolute world coordinate frame. This estimation is performed using the camera projection matrix $H$, ensuring a more robust and coherent 3D pose representation.
@@
$\mathbf{J}=\begin{bmatrix}\frac{\partial x_{w}}{\partial x_{p}}&\frac{\partial x_{w}}{\partial y_{p}}&\frac{\partial x_{w}}{\partial z_{w}}\\ \frac{\partial y_{w}}{\partial x_{p}}&\frac{\partial y_{w}}{\partial y_{p}}&\frac{\partial y_{w}}{\partial z_{w}}\\ 0&0&1\end{bmatrix}.$
@@
Final 3D Pose Estimation
@@
$\arg\max_{\{\hat{\mu}_{P^{i}}\}}$ $\prod_{i}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})^{\top}\Sigma_{P^{i}}^{-1}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})\right)$
$\cdot\prod_{i,j}\exp\left(-\frac{\left(\|\hat{\mu}_{P^{i}}-\hat{\mu}_{P^{j}}\|-l_{ij}\right)^{2}}{2\sigma_{ij}^{2}}\right)$
$\cdot\prod_{k\in\mathcal{L}}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})^{\top}\Sigma_{L^{k}}^{-1}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})\right),$

### v2

@@ Abstract
Autonomous driving systems must operate smoothly in human-populated indoor environments, where challenges arise including limited perception and occlusions when relying only on onboard sensors, as well as the need for socially compliant motion planning that accounts for human psychological comfort zones. These factors complicate accurate recognition of human intentions and the generation of comfortable, socially aware trajectories. To address these challenges, we propose SAP-CoPE, an indoor navigation system that integrates cooperative infrastructure with a novel 3D human pose estimation method and a socially-aware model predictive control (MPC)–based motion planner. In the perception module, an optimization problem is formulated to account for uncertainty propagation in the camera projection matrix while enforcing human joint coherence. The proposed method is adaptable to both single- and multi-camera configurations and can incorporate sparse LiDAR point-cloud data. For motion planning, we integrate a psychology inspired personal-space field using the information from estimated human poses into an MPC framework to enhance socially comfort in human-populated environments. Extensive real-world evaluations demonstrate the effectiveness of the proposed approach in generating socially aware trajectories for autonomous systems.
Note to Practitioners
Indoor autonomous mobile robots operating in public facilities such as hospitals must navigate crowded environments where onboard sensing is often challenged by occlusions and limited fields of view, while also accounting for human comfort in addition to collision avoidance. This paper presents SAP-CoPE, a deployable framework that integrates cooperative infrastructure sensing with socially aware motion planning. Ceiling-mounted sensor nodes provide expanded and occlusion-robust perception through uncertainty-aware 3D human pose estimation from cameras, optionally supplemented by sparse LiDAR data. Leveraging these pose estimates, an MPC-based planner incorporating a human personal space field generates trajectories that balance social comfort, efficiency, and safety while respecting the robot’s kinematic constraints. The proposed architecture is modular and compatible with existing indoor robotic platforms using standard hardware, and real-world experiments demonstrate smoother navigation and improved clearance and travelling time performance compared with conventional baseline methods.
@@ I Introduction
Intelligent indoor mobility systems are gaining increasing attention in healthcare, logistics, and the automotive industry [18, 25, 30]. These systems reduce manual transportation tasks, minimize workplace injuries and improve operational efficiency [17]. In restricted environments, mobile robots primarily focus on avoiding collisions with objects and are programmed to stop when humans are nearby. However, this strategy is less effective in public areas like hospitals and shopping malls, where interactions with people are unavoidable. To enable safe and efficient human-robot interaction, such a system requires real-time obstacle detection, accurate human intention prediction, and a robust motion planning algorithm for safe and smooth navigation.
Recent research has been enhancing the perception ability of a robot through high-end onboard sensors; however, these systems still have limitations in perception range and field of view (FOV), especially in crowded indoor spaces [31]. Ning et al. introduced the Infrastructure Sensor Nodes (ISNs) system, ceiling-mounted multi-modal sensors perform local processing of raw data before transmitting concise perception outputs to a central cloud [23]. This architecture effectively expands the FOV, mitigates occlusions, and offers a cost-efficient alternative to high-end onboard sensing systems. However, accurately estimating pedestrian’s position and heading in the world coordinate using cost-efficient sensors remains a significant challenge. This information is critical to understand human intentions, which in turn enables proactive navigation and effective collision avoidance[37]. To address this, we propose a 3D pose estimation approach using probabilistic fusion of image and sparse point cloud data, ensuring real-time, accurate human pose detection.
Figure 1: Key challenges: (1) How to reliably detect humans in crowded, occluded scenarios? (2) How to guide the robot for human-comfortable motion?
On the other hand, Yang et al. [40] previously introduced a Model Predictive Control (MPC)-based motion planning algorithm that treats all objects as obstacles, prioritizing only physical safety. However, research shows that people maintain social zones in shared spaces [24], highlighting that psychological safety is equally important when navigating interactions with humans. To address this, we integrate a personal space concept from psychology studies into motion planning, enabling robots to respect social behaviour, and maintain comfortable distances in human-populated environments.
@@
    Real-time Cooperative Perception System: We proposed a cooperative perception framework that leverages sensor node networks to provide comprehensive environmental awareness.
@@
    3D Human Pose Detection: We develop a novel approach for 3D human pose detection that combines geometry-based techniques and sensor fusion between image data and sparse point clouds. This method enhances detection accuracy.
@@
    Socially-aware Motion Planning Scheme: We propose an MPC–based motion planning algorithm that integrates the personal space (PS) field to allow the robot to generate socially acceptable motion while accounting for the omnidirectional type robot’s kinematic constraints to ensure the generation of feasible and practical motion of the robot. In addition, the considered PS model has an asymmetrical shape in the lateral direction, which can allow the smooth motion of the robot when encountering a large group of people.
@@
    Validation through Real-world Experiments: The proposed system is extensively validated through real-world experiments. Results demonstrate its accurate pose detection and effectiveness in handling human-robot interactions.
The structure of this article is organized as follows. Section II presents an overview of the proposed methodology. Section III details the perception algorithm. Section IV formulates the socially-aware motion planner. Section V presents the experimental results for the proposed system. Finally, Section VI concludes the study and discusses potential future research directions.
@@ II-A Perception
Infrastructure-Based Perception Systems: Reliable perception is fundamental for intelligent indoor mobility systems. Traditional infrastructure-based approaches, particularly those using vision sensors, have been widely explored. Works such as [42, 13, 11, 7] employ ceiling or wall-mounted cameras to detect and track pedestrians. However, purely independent camera-driven systems can be sensitive to lighting conditions and occlusions, and they are generally limited to obtain precise 3D object positions without additional depth information. To overcome these limitations, researchers have explored multi-sensor fusion within the infrastructure. Brvsvcic et al. [5] integrated RGB-D cameras, LiDAR, and marker-based tracking to enhance spatial accuracy and robustness. Though this approach improves performance, its scalability is hindered by high equipment and installation costs. Similarly, high-frequency motion capture systems [28] offer ground-truth level accuracy but are restricted to controlled spaces, limiting their practical use in real-world deployments. Overall, these limitations emphasize the demand for indoor perception systems that are both reliable under challenging conditions and feasible for cost-effective, large-scale deployment.
Cooperative Perception Systems: Cooperative perception improves environmental awareness by integrating sensory data from multiple sources, including other robots and infrastructure. This collaborative approach effectively addresses challenges such as occlusions and the limited FOV inherent in individual sensors. Recent advancements highlight the significance of Vehicle-to-Everything (V2X) in autonomous driving domain[1, 14, 3, 6]. Huang et al. [14] highlight the critical role of data sharing between agents and infrastructure while identifying real-time performance and communication delay as key bottlenecks. Inspired by this concept, we propose a real-time cooperative perception system utilizing ceiling-mounted ISNs. Each node performs local perception and transmits lightweight observations to a central fusion module in real time. This architecture not only expands FOV and reduces occlusions in crowded indoor spaces, but also ensures timely and coherent perception by mitigating latency-induced errors in the fusion pipeline.
3D Human Pose Detection: Accurately estimating human 3D poses is vital for autonomous systems to interpret and predict human behavior, ensuring safe and socially aware interactions. Single camera-based methods have faced challenges such as depth ambiguity and occlusions[39], thus, the estimation accuracy can be enhanced by leveraging multi-modal sensor data. Zanfir et al. introduced HUM3DIL, a semi-supervised multi-modal approach that embeds LiDAR points into pixel-aligned multi-modal features and employs a Transformer architecture for prediction[41]. Similarly, Bauer et al. proposed a weakly supervised method that integrates camera and LiDAR data, enabling accurate 3D human pose estimation without extensive 3D annotations[4], however, it needs dense point cloud to generate high-accuracy pseudo 3D keypoint labels via projecting point cloud to image.
@@ II-B Planning
Social-aware Planning: Recent human-aware planning algorithms focus primarily on modeling human–robot interactions while often assuming the robot has a simple circular shape for computational convenience, without explicitly accounting for rectangular or articulated platform geometries [8, 9, 34, 33]. Among these algorithms, the Social Force Model (SFM) is widely used to generate socially acceptable robot motion. Kamezaki et al. [15] introduced a planning algorithm based on the inducible Social Force Model, while Ratsamme et al. [27] developed the Extended SFM (ESFM) by incorporating the human pose, orientation, etc. to improve the performance of conventional SFM. Despite their effectiveness, SFM-based approaches often model human influence as symmetric in all directions. While this assumption allows the robot to maintain distance from personal spaces in sparse environments, it can lead to frequent unexpected deceleration or even stopping in crowded settings, thereby reducing efficiency and increasing the risk of collisions.
Planning based on Information Types: Socially-aware planning algorithms can be broadly categorized into two types: individual state-based information and group-based information [38]. Individual state-based methods use detailed information such as position, pose, orientation, and velocity; for example, Stefanini et al. [21] proposed a novel planning algorithm that incorporates the articulated 3D human poses, semantic labels, and trajectory prediction to describe human behaviour. However, such approaches usually require reliable perception and tracking algorithms, and their performance can degrade significantly under occlusions, which are common in cluttered environments.
In contrast, group-based methods treat multiple humans as a single entity. Hoang et al. [22] opposed a proactive social motion model that accounts socio-spatiotemporal characteristics of human groups, allowing the mobile robot to plan a safe trajectory in various situations. Nevertheless, group-level representations may oversimplify complex interactions among individuals, potentially causing the robot to react conservatively or ignore subtle cues such as interpersonal distances, leader–follower structures, or individuals breaking away from the group.
On the other hand, although the above approaches have demonstrated strong performance in human-aware planning, they may not generalize well to large, rectangular robots with omnidirectional maneuverability, potentially resulting in unrealistic or impractical motions. Moreover, these methods often lack explicit constraints on robot kinematics and control limits, which may produce motion commands that are smooth in simulation but difficult to track reliably on actual hardware.
III Methodology
@@
The proposed SAP-CoPE method consists of two main systems: a real-time cooperative perception and a socially aware motion planning (Fig. 2). For the perception systems, it employs a group of multi-modal ISNs equipped with LiDAR, cameras, and edge computing for local object detection and 3D human pose estimation. A synchronized clock system aligns LiDAR scans with camera operations, ensuring temporally consistent data, which is transmitted to a Central Layer (CL) for global fusion, multi-object tracking, and delay compensation. For the motion-planning system, an MPC-based planner is designed to generate optimal poses and is implemented in the CL, enabling the robot to compute a socially-aware motion that respects humans’ personal space while close to a predefined reference motion. To achieve this, the PS field inspired from psycholog studies is incorporated into the MPC’s objective function.
Furthermore, the optimal control commands can then be obtained from the controller by tracking the motion generated from the planner. And these commands are then transmitted to the Application Layer (AL) for execution. In this study, the communication between the ISNs, CL, and AL is achieved using Wi-Fi network.
Additionally, the proposed SAP-CoPE framework is adaptable to various robotic platforms, such as the autonomous cart and autonomous medical bed at the Mechatronic Vehicle Systems Lab, University of Waterloo. In this study, only the autonomous medical bed [40] was used for experimental validation and is referred to as the term ”robot” throughout this paper.
@@ IV-A1 Camera-based 2D Detection
Recent advancements in deep learning have significantly improved camera-based 2D object detection. Notably, YOLOv11 [16] demonstrates superior accuracy and inference speed due to its state-of-the-art backbone and neck architectures. However, the pre-trained model does not generalize well to the top-down sensor node view and is unable to detect the autonomous robot used in our application. To address these limitations, we fine-tuned the YOLOv11 model using a dataset combining COCO [19] and the Indoor Cooperative Infrastructure Dataset (ICID) [23] to achieve human 2D pose detection and robot bounding box detection in one single network.
@@ IV-A2 LiDAR Processing and Class Association
The LiDAR processing consists of region-of-interest filtering and a hierarchical clustering algorithm to group the points for obstacles while considering the scanning pattern. The clustering results are then fused with the camera-based 2D detection results for robot and human pose estimation. More details on the implementation of clustering and association can be found in our previous work [23]. The novel proposed 3D human pose estimation method is discussed in the following section.
@@
Estimating absolute 3D human pose from a single camera is challenging due to the lack of depth information. As indicated in Fig. 2, our approach first obtains the 2D pixel coordinates of each joint and leverages prior height distribution information to estimate the rough 3D distribution of each joint as $P^{i}\sim N(\mu_{P^{i}},\Sigma_{P^{i}})$ in the absolute world coordinate frame. This estimation is performed using the camera projection matrix $H$, ensuring a more robust and coherent 3D pose representation.
@@
$\mathbf{J}=\begin{bmatrix}\frac{\partial x_{w}}{\partial x_{p}}&\frac{\partial x_{w}}{\partial y_{p}}&\frac{\partial x_{w}}{\partial z_{w}}\\ \frac{\partial y_{w}}{\partial x_{p}}&\frac{\partial y_{w}}{\partial y_{p}}&\frac{\partial y_{w}}{\partial z_{w}}\\ 0&0&1\end{bmatrix}$ (5)
@@
Final 3D Pose Estimation (Node Level)
@@
$\arg\max_{\{\hat{\mu}_{P^{i}}\}}\$$\prod_{i}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})^{\top}\Sigma_{P^{i}}^{-1}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})\right)$
$\cdot\prod_{i,j}\exp\left(-\frac{\left(\|\hat{\mu}_{P^{i}}-\hat{\mu}_{P^{j}}\|-l_{ij}\right)^{2}}{2\sigma_{ij}^{2}}\right)$ (6)
$\cdot\prod_{k\in\mathcal{L}}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})^{\top}\Sigma_{L^{k}}^{-1}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})\right).$